Solve for x.
sqrt(3x-2) = x.
Can x be 1 or 2. It only makes sense to me as 1.

Answer by scott8148(6628)

both

squaring ___ 3x - 2 = x^2

0 = x^2 - 3x + 2

factoring ___ 0 = (x - 1)(x - 2)

x - 1 = 0 ___ x = 1 ___ sqrt[(3 * 1) - 2] = sqrt(1) = 1

x - 2 = 0 ___ x = 2 ___ sqrt[(3 * 2) - 2] = sqrt(4) = 2